OpenStudy (anonymous):

For a complex number written in polar form, its representation is unique as long as the argument is restricted to (0,2 pi) A. True B. False

OpenStudy (paxpolaris):

\[\Large 0 \le \theta < 2\pi\] as long as you include 0 or 2π but not both